You have 6 red balls, three of which are identical and the other three are distinct and different from the previous three, 4 distinct yellow balls, 5 identical blue balls. You want to select 7 balls to make a gift bag. How many different options per gift bag do you have?
Answer #1

Sol:

Total balls = 6 + 4 + 5 = 15

Balls in which 3 red balls are identical and 5 blue balls are identical.

Total ways to select 7 balls = 15P7 / (3!*5!) = 45045